the rectangle has a premiter of 64cm.
the length needs 4cm. to be twice the size as the width
please slove and explain

Answer by ptaylor(2198)

Perimeter of rectangle=2*length+2*width=2L+2W=64 cm
Let x=width
Then 2x-4=length
2L+2W=2x+2(2x-4)=64cm
2x+2(2x-4)=64 cm get rid of parens
2x+4x-8=64 cm add 8 to both sides
2x+4x-8+8=64+8 collect like terms
6x=72 divide both sides by 6
x=12 cm---------------width
2x-4=2*12-4=20 cm-------------length
CK
2L+2W=40+24=64cm
Hope this helps----ptaylor